OpenStudy (anonymous):

Right circular cone the height= diameter when height = 2.98 cm find the volume of the cone (by approximation using mean value theorem )

OpenStudy (anonymous):

tell me how i start or give me hints and i will complete

OpenStudy (anonymous):

lol uhm circle is pie radius squared

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

@kadkins15 i know the volume of cone \[V=\frac{ 1 }{ 3 }\pi r ^{2}h\]

OpenStudy (anonymous):

okay well i k that h is 2.98 pie is 3.14 and your radius^3 is what ur trying to find

OpenStudy (anonymous):

trying to find the volume by approximation using mean value theorem

OpenStudy (anonymous):

@ganeshie8 can u plz help me with this

OpenStudy (anonymous):

v=1/2(3.14)r^3(2.8)

OpenStudy (anonymous):

i meant 1/2(3.14)

OpenStudy (anonymous):

1/3(3.14

OpenStudy (anonymous):

i know it's easy to find it using calculator by i want to find it using approximation by mean value theorem

OpenStudy (anonymous):

give me a second ill b rite bck

OpenStudy (fantalux):

http://www.sosmath.com/calculus/diff/der11/der11.html Does this help at all? XD

OpenStudy (anonymous):

@Fantalux i know everything about MVT it's about the question , no it doesn't help thanks anyway lol

OpenStudy (anonymous):

is this calculus or what??

OpenStudy (fantalux):

ikr

OpenStudy (anonymous):

no the teacher just game the question i don't know from where did he bring it

OpenStudy (anonymous):

gave *

OpenStudy (anonymous):

What does this mean? As usual, let us utilize an example to grasp the concept. Visualize (or graph) the function f(x)=x^{3}. Choose an interval (anything will work), but for the sake of simplicity, [0,2]. Draw a line going from point (0,0) to (2,8). Between the points x = 0 and x = 2 exists a number x = c, where the derivative of f at point c is equal to the slope of the line you drew. Solution: 1: Using the definition of the mean value theorem {f(b)-f(a) \over b-a} insert values. Our chosen interval is [0,2]. So, we have {\frac {f(2)-f(0)}{2-0}}={\frac {8}{2}}=4 2: By the definition of the mean value theorem, we know that somewhere in the interval exists a point that has the same slope as that point. Thus, let us take the derivative to find this point x = c. {\frac {dy}{dx}}=3x^{2} Now, we know that the slope of the point is 4. So, the derivative at this point c is 4. Thus, 4=3x^{2}. The square root of 4/3 is the point.

OpenStudy (anonymous):

does that help at all

OpenStudy (ikram002p):



OpenStudy (ikram002p):

volume = 1/3 ((0.5h)^2 pi))*h

OpenStudy (anonymous):

@ikram002p i know how to find the area of cone but i want to use approximation using mean value theorem

OpenStudy (ikram002p):



OpenStudy (ikram002p):

lol i knw .. just find a function of volume then use mean value thm

OpenStudy (anonymous):

ok :)

OpenStudy (anonymous):

hope that helps i gotta go bro but yeahh

OpenStudy (anonymous):

thanks bro

OpenStudy (ikram002p):

so u have h=2.98 .... h'=3 v=pi/12 h^3 v'=pi/4 h^2 \[f(h)\approx f(h')+f'(h')(h-h')\]

OpenStudy (anonymous):

aha thanks @ikram002p

OpenStudy (anonymous):

i've made a mistake in the value of h' when i solved it first time lol

OpenStudy (ikram002p):

so its good nw ha ?!

OpenStudy (anonymous):

yup thank you so much :)

OpenStudy (ikram002p):

np :)